Transverse waves with a speed of 42.0 m/s are to be produced on a stretched string. A 4.60-m length of string with a total mass of 0.0600 kg is used.

(a) What is the required tension in the string?

. (b) Calculate the wave speed in the string if the tension is 8.00 N.
Answer #1

a) let the tension in the string is T

as v = sqrt(T/u)

42 = sqrt(T * 4.6/0.006)

solving for T

T = 2.301 N

the tension in the string is 2.301 N

b)

Now ,

v1/v2 = sqrt(T1/T2)

v1/42 = sqrt(8/2.3)

solving for v1

v1 = 78.33 m/s

the wave speed of the string is 78.33 m/s